# Geodesic Slice Sampling on the Sphere

### v4

@@ 1 Introduction
    Posteriors naturally defined on the sphere require efficient sampling: To provide some examples, sampling distributions on the sphere play an important role in directional statistics and shape analysis (see Mardia and Jupp, 2000), and Bayesian inverse problems on $\mathbb{S}^{2}$ occur, e.g., in astrophysics or geophysics (see Marignier et al., 2023). Moreover, Bayesian density estimation (Holbrook et al., 2020) requires sampling of a posterior on an infinite-dimensional sphere, which usually is approximated by truncating the dimension, ending up with $\mathbb{S}^{d-1}$ for large $d$. Also the problem of 3D rigid registration can be reduced to sampling from a distribution on $\mathbb{S}^{3}$ (see Section 4.1).
@@
Let us comment on how our contributions fit into the literature. Lie et al. (2023) present two MCMC algorithms to sample distributions on the sphere using push forward kernels of a preconditioned Crank-Nicolson algorithm and of elliptical slice sampling, respectively. The crucial difference to our work is that Lie et al. (2023) assume that the target density is defined relative to an angular Gaussian distribution. Instead, here we consider target densities relative to the volume measure of the sphere. Observe that with increasing dimension both reference measures become “increasingly singular” to each other. Furthermore, Marignier et al. (2023) propose a proximal MCMC method to sample from posterior distributions of inverse imaging problems on $\mathbb{S}^{2}$. An infinite-dimensional setting of Bayesian density estimation is treated by Holbrook et al. (2020) by an HMC algorithm on a sphere. Other MCMC approaches have been developed on more general manifolds that usually cover $\mathbb{S}^{d-1}$ as a special case. This includes the HMC algorithm for manifolds embedded in the Euclidean space introduced by Byrne and Girolami (2013) which uses the geodesic flow, and the HMC algorithm for manifolds that are the fibre of a smooth map proposed by Brubaker et al. (2012). Furthermore, Goyal and Shetty (2019) investigate a Metropolis-Hastings-like geodesic walk on manifolds with non-negative curvature. Similarly, the Metropolis-Hastings algorithm of Zappa et al. (2018) based on projections from the tangent spaces can be applied to $\mathbb{S}^{d-1}$. Here it is also worth to mention that several geodesic walks exist that target the uniform distribution on subsets of the sphere. We refer to Section 2.3 for more details. Finally, note that recently there has been some theoretical progress in the investigation of convergence properties of slice sampling (see for example Łatuszyński and Rudolf, 2024; Natarovskii et al., 2021b, a; Hasenpflug et al., 2024) that very much influenced the presentation and proof arguments of our theoretical results.
@@ 2.2 The Geometry of the Sphere
We continue by illuminating an interaction between geometric and measure theoretic structure of the sphere with the help of the map
@@
2.3 Geodesic Walk on the Sphere.
@@
Related work.
If the step-size is chosen to be constant, that is, $\tau=\delta_{\varepsilon}$ for $\delta_{\varepsilon}$ being the Dirac measure on $\mathbb{R}$ at $\varepsilon\in(0,2\uppi)$, then $K_{\delta_{\varepsilon}}$ coincides with the geodesic walk from Mangoubi and Smith (2018) on the sphere. They provide dimension independent mixing time results for this walk on manifolds with bounded positive sectional curvature.
@@ 3.1 Ideal Geodesic Slice Sampling
Observe that the size of $r_{p}\in(0,1]$ solely depends on the decreasing level set function $t\mapsto\mathcal{U}_{\mathbb{S}^{d-1}}(L(t))$. This function plays an even stronger role in controlling the convergence of ideal slice sampling. It is known to be the only relevant quantity for geometric ergodicity and spectral gap estimates, cf. Roberts and Rosenthal (1999); Natarovskii et al. (2021b).
@@
We can interpret the ideal geodesic slice sampler as hit-and-run on the sphere. The hit-and-run algorithm has been introduced for approximate sampling of distributions on $\mathbb{R}^{d}$ (see Bélisle et al., 1993). A transition of the algorithm that starts in $x\in\mathbb{R}^{d}$ is performed in the following way: A direction $v$ is chosen uniformly at random in the unit ball centered around $x$ to construct a straight line $\ell$ through $x$ in direction $v$. Then the next point is chosen w.r.t. the distribution of interest restricted to the line $\ell$. Since on the sphere straight lines correspond to great circles, ideal geodesic slice sampling can be considered as hit-and-run where simulating the target distribution restricted to the great circle is performed by slice sampling. Investigations on the performance of the hit-and-run algorithm in $\mathbb{R}^{d}$ w.r.t. the dimension attracted a lot of attention, cf. Lovász and Vempala (2007). Regarding this aspect note that for $\varepsilon\in(0,1)$ we require $n\geq\sqrt{2(d-1)\uppi}\,r_{p}^{-1}\log\varepsilon^{-1}$ for achieving $\sup_{x\in\mathbb{S}^{d-1}}d_{tv}\left(H^{n}(x,\cdot),\pi\right)<\varepsilon$, which illuminates a moderate dependence on $d$.
@@
$\sup_{x\in\mathbb{S}^{d-1}}d_{tv}\left(P^{n}(x,\cdot),\eta\right)\leq(1-\nu(\mathbb{S}^{d-1}))^{n},\qquad\forall n\in\mathbb{N}.$
@@
$\int_{0}^{\uppi}\int_{{\mathbb{S}^{d-2}_{x}}}\mathbbm{1}_{A}(\gamma_{(x,v)}(\theta))\,\mathcal{U}_{{\mathbb{S}^{d-2}_{x}}}({\rm d}v)\,{\rm d}\theta\geq\frac{\sqrt{2\uppi}}{\sqrt{d-1}}\;\mathcal{U}_{\mathbb{S}^{d-1}}(A).$
@@
3.2 Geodesic Shrinkage Slice Sampling.
@@
The distribution $Q_{{x,}{v,}{t}}$ coincides with the kernel of the shrinkage procedure defined by Hasenpflug et al. (2024, Algorithm 2.2 with $\theta_{\text{in}}=0$, $S=L(x,v,t)$). For details we refer to Appendix E. The lower semicontinuity of $p$ ensures that $Q_{{x,}{v,}{t}}$ is well-defined, cf. Hasenpflug et al. (2024). Intuitively, the openness of the superlevel sets guarantees that $x$ lies in a subinterval of the image of the geodesic level set $\gamma_{(x.v)}\big{(}L(x,v,t)\big{)}$ for all $v\in{\mathbb{S}^{d-2}_{x}}$ and all $t\in(0,p(x))$. Hence, while shrinking the sampling region on $\gamma_{(x,v)}\big{(}[0,2\uppi)\big{)}$, the probability to hit the level set for a given level $t\in(0,p(x))$ is strictly positive.
@@
Proof (Proof of Theorem 16.) For arbitrary $x\in\mathbb{S}^{d-1}$, $s\in(0,\left\|p\right\|_{\infty})$, for any $v\in{\mathbb{S}^{d-2}_{x}}$ and $\theta\in[0,2\uppi)$ with $\gamma_{(x,v)}(\theta)\in L(s)$ we have (cf. (8)) that
@@
$Q_{{x,}{v,}{t}}(B)\geq\mathbb{P}(\Xi\in B\cap L(x,v,t))=\frac{1}{2\uppi}{\lambda}(B\cap L(x,v,t)),\quad B\in\mathcal{B}([0,2\uppi)).$ (11)
@@
Note that the right hand-side of (9) coincides with the the right hand-side of (7) stated in Theorem 6 that addressed the uniform ergodicity statement for ideal slice sampling. This is due to the proof technique, since in the small set estimate, we lower bound both kernels with the same expression appearing in Lemma 11. Intuitively, it is clear that $\sup_{x\in\mathbb{S}^{d-1}}d_{tv}(H^{n}(x,\cdot),\pi)$ is smaller than $\sup_{x\in\mathbb{S}^{d-1}}d_{tv}(\widetilde{H}^{n}(x,\cdot),\pi)$, since the shrinkage procedure just adaptively imitates the acceptance/rejection step of the ideal one to gain computational efficiency. Exactly this gain in efficiency leads in applications to a potentially better accuracy to cost ratio, although the performance per Markov chain transition may be worse. We also point out that Remark 7 and Remark 8 apply to Theorem 16 as well.
@@ 4.1 Rigid Registration of Biomolecular Structures
$p(\boldsymbol{q}_{i}\mid P,\boldsymbol{R},\omega,\sigma)=\omega p_{o}(\boldsymbol{q}_{i})+\frac{1-\omega}{J\,(2\pi\sigma^{2})^{3/2}}\sum_{j=1}^{J}\exp\left(-\frac{1}{2\sigma^{2}}\|\boldsymbol{q}_{i}-\boldsymbol{R}\boldsymbol{p}_{j}\|^{2}\right)\,.$ (12)
@@
In the following, we assume that the fraction of outliers $\omega$ and the standard deviation $\sigma$ are known such that the only unknown parameter is the unit quaternion $x$ encoding a rotation of the source relative to the target. We consider an example from structural bioinformatics and test MCMC samplers to tackle this rigid registration task. Adenylate kinase (AK) undergoes a large conformational change upon ligand binding. This internal structural change cannot be modeled as a global rigid transformation. Only the core domain, comprising a 60% subset of all points, can be related to each other via a rigid transformation. In our tests, the target and source point clouds show the closed and open conformation of AK (the protein data bank (PDB) codes 1AKE and 4AKE; see Fig. 3). We fix $\sigma=1$Å and $\omega=0.4$ (motivated by the fact that the core domain contains roughly 60% of all points); the bounding box volume of 1AKE is $5.73\times 10^{4}{\rm\AA }^{3}$. To guide the visual inspection, we use the chainbow coloring from blue to red offered by the Pymol software. Note that the posterior (14) is invariant under reordering of the points.
@@
The posterior density (14) has many local peaks as is clear from Figure 4A showing a slice through $p$ along a random great circle passing through the posterior mode. Rigid registration by sampling from posterior (14) is challenging, because standard MCMC samplers tend to get stuck in subordinate peaks and miss the global maximum. To gain an overview of the posterior, we systematically evaluate $p$ over a regular tessellation of $\mathbb{S}^{3}$ based on the 600-cell (Straub et al., 2017). The 600-cell is the 4D analog of the regular icosahedron and can be used to approximate $\mathbb{S}^{3}$ at varying degrees of resolution. A single cell is a 3D tetrahedron whose corners lie in $\mathbb{S}^{3}$. By splitting each tetrahedron into eight equally-sized tetrahedra we obtain a tessellation at a higher resolution. Starting from the 600-cell encoding 300 unique rotations, we repeat the splitting process four times to obtain 1279264 unit quaternions that tessellate $\mathbb{S}^{3}$ in a regular fashion. Evaluation of $p$ at these quaternions offers an accurate discretization of the posterior that allows us to study its shape (see Figure 4B). The global maximum of $\log p$ on the quaternions of the 600-cell tessellation is 2192.89-2192.89. Although, an overwhelming fraction of quaternions achieves only very small posterior probabilities, almost all probability mass is concentrated in the maximum posterior peak: the fractional volume of $\mathbb{S}^{3}$ covered by quaternions $x\in\mathbb{S}^{3}$ with $\log p(x)>-2300$ is close to zero ($1.27\times 10^{-4}$ based on the 600-cell tessellation), but the cumulative probability is close to one ($1-4\times 10^{-44}$).
@@ 4.2 Mixture of von Mises-Fisher Distributions
where every $\mu_{k}\in\mathbb{S}^{d-1}$ with $k\in\{1,\dots,K\}$ is sampled w.r.t. the uniform distribution on the unit sphere and then fixed. We evaluate both geoSSS variants against RWMH and HMC on a $10$-dimensional mixture of vMF distributions with $K=5$ mixture components on the unit sphere $\mathbb{S}^{9}$. We set $\kappa=100$ and draw $10^{6}$ samples.
@@
To gain some insight into the performance of the samplers on the multi-modal target, we first look at the marginal distributions (see Fig. 5). We estimate marginal histograms from the samples and compare them with the marginal distributions of the vMF mixture that we computed analytically. Both variants of geoSSS represent the underlying mixture significantly better compared to HMC and especially RWMH which seems to explore only a single component of the vMF mixture.
Next we study quality measures to obtain more quantitative insights into the sampling performance. RWMH is the fastest to decorrelate, which is reflected by the rapid decay of the autocorrelation function (ACF) (see Fig. 6A). However, the reason for the fast decorrelation is that RWMH fails to explore the entire distribution, but is stuck in a single component. The geoSSS variants decorrelate much faster as compared to HMC. Note that the slow, linear decay of the ACF results from the fact that HMC spends a large fraction of the simulation time in a single mode and only occasionally escapes to another mode (also HMC only visits three out of all five modes in the course of the simulation).
@@
where $\ell=(\ell_{1},\ldots,\ell_{K})$ with $\ell_{k}=1/K$ for $k\in\{1,\ldots,K\}$ is the ideal distribution and $q=(q_{1},\ldots,q_{k})$. To estimate $q_{k}$, we assign each sample $x_{n}$ to the component under which $x_{n}$ is most likely, i.e.,
@@
Based on the allocations $k_{n}$, we can count how often each mode is visited by the Markov chain. A good MCMC sampler should produce $q_{k}\approx 1/K$ resulting in $\text{KL}(q\mid\ell)\approx 0$. The more the KL divergence differs from the minimum value of zero, the greater is the mismatch between the ideal and empirical frequency of mode visits. As seen from Fig. 6B, only the geoSSS variants achieve a small KL divergence. RWMH and HMC produce large KL divergences, indicating a complete failure to detect some of the components and, as a result, a poorer representation of the target distribution compared to geoSSS. Indeed, RWMH is stuck in a single mode in the course of the simulation. HMC fares better, but also misses two out of five modes.
@@
An efficient MCMC algorithm should explore the sphere rapidly by making large leaps from one sample to the next. The great circle distance between successive samples (see Fig. 6C) shows that the slice samplers explore the sphere more efficiently than HMC and RWMH.
We further evaluate the performance of the samplers by fixing the components on the sphere and increasing the concentration parameter $\kappa$ from $50$ to $500$, therefore making the distributions “spikier”. We estimate the effective sample size (ESS) for each method by considering $10$ chains and $10^{6}$ MCMC steps from the first dimension, resulting in ESS values per method for each $\kappa$. Figure 7A shows the estimated ESS for each of the four samplers as a function of the concentration parameter $\kappa$. Overall, we observe that sampling the mixture model becomes more challenging as the components become more concentrated with increasing $\kappa$ for all methods. However, both geoSSS outperform RWMH and HMC, particularly for lower $\kappa$ values. At large $\kappa$ values, all samplers perform similarly poorly and fail to produce a reliable approximation of the target.
@@
It is worth noting that the rejection-based geoSSS yields the highest ESS, albeit at the cost of generating a significantly larger number of rejected samples in comparison to the shrinkage-based geoSSS. This is demonstrated in Fig. 7B, where it is observed that the rejection-based geoSSS generates an increasingly larger number of rejected samples for higher values of $\kappa$. For instance, for $\kappa=50$, approximately 17 rejections happen per MCMC step, while for $\kappa=500$, around 60 rejections occur. On the other hand, the number of rejections for the shrinkage-based geoSSS increases at a much slower rate with $\kappa$. It can be seen here that for $\kappa=50$, we have approximately 4 rejections per MCMC step, and for $\kappa=500$, approximately just 6 rejections per MCMC step occur. The number of rejections directly determines the number of log probability evaluations and thereby the computational costs of the slice samplers. RWMH and HMC, on the other hand, have a fixed computational budget per MCMC step. In case of RWMH, one step requires a single log probability evaluation, whereas an HMC step involves one log probability evaluation as well as multiple gradient evaluations, one for each leapfrog step. In our settings, we use 10 leapfrog integration steps. In summary, RWMH has the lowest computational costs, but also the smallest ESS. The computational costs of shrinkage-based geoSSS is smaller than the costs of HMC, whilst it also performs better than HMC in terms of the other evaluation criteria.

### v5

@@ 1 Introduction
    Posteriors naturally defined on the sphere require efficient sampling: To provide some examples, sampling distributions on the sphere plays an important role in directional statistics and shape analysis (see Mardia and Jupp, 2000), and Bayesian inverse problems on $\mathbb{S}^{2}$ occur, e.g., in astrophysics or geophysics (see Marignier et al., 2023). Moreover, Bayesian density estimation (Holbrook et al., 2020) requires sampling of a posterior on an infinite-dimensional sphere, which usually is approximated by truncating the dimension, ending up with $\mathbb{S}^{d-1}$ for large $d$. Also the problem of 3D rigid registration can be reduced to sampling from a distribution on $\mathbb{S}^{3}$ (see Section 4.1).
@@
Let us comment on how our contributions fit into the literature. Lie et al. (2023) present two MCMC algorithms to sample distributions on the sphere using push forward kernels of a preconditioned Crank-Nicolson algorithm and of elliptical slice sampling, respectively. The crucial difference to our work is that Lie et al. (2023) assume that the target density is defined relative to an angular Gaussian distribution. Instead, here we consider target densities relative to the volume measure of the sphere. Observe that with increasing dimension both reference measures become “increasingly singular” to each other. Furthermore, Marignier et al. (2023) propose a proximal MCMC method to sample from posterior distributions of inverse imaging problems on $\mathbb{S}^{2}$. An infinite-dimensional setting of Bayesian density estimation is treated by Holbrook et al. (2020) by an HMC algorithm on a sphere. Other MCMC approaches have been developed on more general manifolds that usually cover $\mathbb{S}^{d-1}$ as a special case. This includes the HMC algorithm for manifolds embedded in the Euclidean space introduced by Byrne and Girolami (2013) which uses the geodesic flow, and the HMC algorithm for manifolds that are the fibre of a smooth map proposed by Brubaker et al. (2012). Furthermore, Goyal and Shetty (2019) investigate a Metropolis-Hastings-like geodesic walk on manifolds with non-negative curvature. Similarly, the Metropolis-Hastings algorithm of Zappa et al. (2018) based on projections from the tangent spaces can be applied to $\mathbb{S}^{d-1}$. Here it is also worth to mention that several geodesic walks exist that target the uniform distribution on subsets of the sphere. We refer to Section 2.3 for more details. Finally, note that recently there has been some theoretical progress in the investigation of convergence properties of slice sampling (see for example Natarovskii et al., 2021a, b; Łatuszyński and Rudolf, 2024; Hasenpflug et al., 2024) that very much influenced the presentation and proof arguments of our theoretical results.
@@ 2.2 The Geometry of the Sphere
We continue by illuminating an interaction between the geometric and the measure theoretic structure of the sphere with the help of the map
@@
2.3 Geodesic Walk on the Sphere
@@
Related work. If the step-size is chosen to be constant, that is, $\tau=\delta_{\varepsilon}$ for $\delta_{\varepsilon}$ being the Dirac measure on $\mathbb{R}$ at $\varepsilon\in(0,2\uppi)$, then $K_{\delta_{\varepsilon}}$ coincides with the geodesic walk from Mangoubi and Smith (2018) on the sphere. They provide dimension independent mixing time results for this walk on manifolds with bounded positive sectional curvature.
@@ 3.1 Ideal Geodesic Slice Sampling
Observe that for lower semicontinuous $p$ the size of $r_{p}\in(0,1]$ solely depends on the decreasing level set function $t\mapsto\mathcal{U}_{\mathbb{S}^{d-1}}(L(t))$. This function plays an even stronger role in controlling the convergence of ideal slice sampling. It is known to be the only relevant quantity for geometric ergodicity and spectral gap estimates, cf. Roberts and Rosenthal (1999); Natarovskii et al. (2021b).
@@
We can interpret the ideal geodesic slice sampler as hit-and-run on the sphere. The hit-and-run algorithm has been introduced for approximate sampling of distributions on $\mathbb{R}^{d}$ (see Bélisle et al., 1993). A transition of the algorithm that starts in $x\in\mathbb{R}^{d}$ is performed in the following way: A direction $v$ is chosen uniformly at random in the unit ball centered around $x$ to construct a straight line $\ell$ through $x$ in direction $v$. Then the next point is chosen w.r.t. the distribution of interest restricted to the line $\ell$. Since on the sphere straight lines correspond to great circles, ideal geodesic slice sampling can be considered as hit-and-run where simulating the target distribution restricted to the great circle is performed by slice sampling. Investigations on the performance of the hit-and-run algorithm in $\mathbb{R}^{d}$ w.r.t. the dimension attracted a lot of attention, cf. Lovász and Vempala (2007). Regarding this aspect note that for $\varepsilon\in(0,1)$ we require $n\geq\sqrt{2(d-1)\uppi}\,r_{p}^{-1}\log(\varepsilon^{-1})$ for achieving $\sup_{x\in\mathbb{S}^{d-1}}d_{tv}\left(H^{n}(x,\cdot),\pi\right)<\varepsilon$, which illuminates a moderate dependence on $d$.
@@
$\sup_{x\in\mathbb{S}^{d-1}}d_{tv}\left(P^{n}(x,\cdot),\eta\right)\leq\big(1-\nu(\mathbb{S}^{d-1})\big)^{n},\qquad\forall n\in\mathbb{N}.$
@@
$\int_{0}^{\uppi}\int_{{\mathbb{S}^{d-2}_{x}}}\mathbbm{1}_{A}\left(\gamma_{(x,v)}(\theta)\right)\,\mathcal{U}_{{\mathbb{S}^{d-2}_{x}}}({\rm d}v)\,{\rm d}\theta\geq\frac{\sqrt{2\uppi}}{\sqrt{d-1}}\;\mathcal{U}_{\mathbb{S}^{d-1}}(A).$
@@
3.2 Geodesic Shrinkage Slice Sampling
@@
The distribution $Q_{{x,}{v,}{t}}$ coincides with the kernel of the shrinkage procedure defined by Hasenpflug et al. (2024, Algorithm 2.2 with $\theta_{\text{in}}=0$, $S=L(x,v,t)$). For details we refer to Appendix E. The lower semicontinuity of $p$ ensures that $Q_{{x,}{v,}{t}}$ is well-defined, cf. Hasenpflug et al. (2024). Intuitively, the openness of the superlevel sets guarantees that $x$ lies in a subinterval of the image of the geodesic level set $\gamma_{(x.v)}\big(L(x,v,t)\big)$ for all $v\in{\mathbb{S}^{d-2}_{x}}$ and all $t\in(0,p(x))$. Hence, while shrinking the sampling region on $\gamma_{(x,v)}([0,2\uppi))$, the probability to hit the level set for a given level $t\in(0,p(x))$ is strictly positive.
@@
Proof (Proof of Theorem 16.) For arbitrary $x\in\mathbb{S}^{d-1}$, $s\in(0,\left\|p\right\|_{\infty})$, for any $v\in{\mathbb{S}^{d-2}_{x}}$ and $\theta\in[0,2\uppi)$ with $\gamma_{(x,v)}(\theta)\in L(s)$ we have (cf. Equation 8) that
@@
$Q_{{x,}{v,}{t}}(B)\geq\mathbb{P}\big(\Xi\in B\cap L(x,v,t)\big)=\frac{1}{2\uppi}{\lambda}\big(B\cap L(x,v,t)\big),\quad B\in\mathcal{B}([0,2\uppi)).$ (11)
@@
Note that the right hand-side of (9) coincides with the the right hand-side of (7) stated in Theorem 6, that addressed the uniform ergodicity statement for ideal slice sampling. This is due to the proof technique, since in the small set estimate, we lower bound both kernels with the same expression appearing in Lemma 11. Intuitively, it is clear that $\sup_{x\in\mathbb{S}^{d-1}}d_{tv}(H^{n}(x,\cdot),\pi)$ is smaller than $\sup_{x\in\mathbb{S}^{d-1}}d_{tv}(\widetilde{H}^{n}(x,\cdot),\pi)$, since the shrinkage procedure just adaptively imitates the acceptance/rejection step of the ideal one to gain computational efficiency. Exactly this gain in efficiency leads in applications to a potentially better accuracy to cost ratio, although the performance per Markov chain transition may be worse. We also point out that Remark 7 and Remark 8 apply to Theorem 16 as well.
@@ 4.1 Rigid Registration of Biomolecular Structures
$p(\boldsymbol{q}_{i}\mid P,\boldsymbol{R},\omega,\sigma)=\omega p_{o}(\boldsymbol{q}_{i})+\frac{1-\omega}{J\,(2\uppi\sigma^{2})^{3/2}}\sum_{j=1}^{J}\exp\left(-\frac{1}{2\sigma^{2}}\|\boldsymbol{q}_{i}-\boldsymbol{R}\boldsymbol{p}_{j}\|^{2}\right)\,.$ (12)
@@
In the following, we assume that the fraction of outliers $\omega$ and the standard deviation $\sigma$ are known such that the only unknown parameter is the unit quaternion $x$ encoding a rotation of the source relative to the target. We consider an example from structural bioinformatics and test MCMC samplers to tackle this rigid registration task. Adenylate kinase (AK) undergoes a large conformational change upon ligand binding. This internal structural change cannot be modeled as a global rigid transformation. Only the core domain, comprising a 60% subset of all points, can be related to each other via a rigid transformation. In our tests, the target and source point clouds show the closed and open conformation of AK (the protein data bank (PDB) codes 1AKE and 4AKE; see Fig. 3). We fix $\sigma=1$Å and $\omega=0.4$ (motivated by the fact that the core domain contains roughly 60% of all points); the bounding box volume of 1AKE is $5.73\times 10^{4}$Å3. To guide the visual inspection, we use the chainbow coloring from blue to red offered by the Pymol software. Note that the posterior (14) is invariant under reordering of the points.
@@
The posterior density (14) has many local peaks as is clear from Figure 4A showing a slice through $p$ along a random great circle passing through the posterior mode. Rigid registration by sampling from posterior (14) is challenging, because standard MCMC samplers tend to get stuck in subordinate peaks and miss the global maximum. To gain an overview of the posterior, we systematically evaluate $p$ over a regular tessellation of $\mathbb{S}^{3}$ based on the 600-cell (Straub et al., 2017). The 600-cell is the 4D analog of the regular icosahedron and can be used to approximate $\mathbb{S}^{3}$ at varying degrees of resolution. A single cell is a 3D tetrahedron whose corners lie in $\mathbb{S}^{3}$. By splitting each tetrahedron into eight equally-sized tetrahedra we obtain a tessellation at a higher resolution. Starting from the 600-cell encoding 300 unique rotations, we repeat the splitting process four times to obtain 1279264 unit quaternions that tessellate $\mathbb{S}^{3}$ in a regular fashion. Evaluation of $p$ at these quaternions offers an accurate discretization of the posterior that allows us to study its shape (see Figure 4B). The global maximum of $\log p$ on the quaternions of the 600-cell tessellation is 2192.89-2192.89. Although, an overwhelming fraction of quaternions achieves only very small posterior probabilities, almost all probability mass is concentrated in the maximum posterior peak: the fractional volume of $\mathbb{S}^{3}$ covered by quaternions $x\in\mathbb{S}^{3}$ with $\log p(x)>-2300$ is close to zero ($1.27\times 10^{-4}$ based on the 600-cell tessellation), but the cumulative probability is close to one ($1-4\times 10^{-44}$). In hindsight, due to the high concentration of the target distribution one could argue that the ability to identify the mode, effectively serving as an optimizer, is of primary importance, while posterior-based uncertainty quantification is less relevant.
@@ 4.2 Mixture of von Mises-Fisher Distributions
where every $\mu_{k}\in\mathbb{S}^{d-1}$ with $k\in\{1,\dots,K\}$ is sampled w.r.t. the uniform distribution on the unit sphere and then fixed.
For this target distribution we also consider a Metropolis algorithm with mixture proposal based on a random walk and independent sampling, which is referred to as mixture-MH. The random walk part allows local exploration, whereas the independence sampling, regarding the uniform distribution on the sphere, encourages global moves. In this scheme the mixture hyperparameter $\alpha\in[0,1]$ appears: With probability $\alpha$ a random walk and with probability $1-\alpha$ an independent sampling proposal is used. The resulting Metropolis kernel can also be interpreted as a mixture of RWMH and independent Metropolis. For implementation details we refer to Algorithm 8 under Appendix F.
We evaluate both geoSSS variants against RWMH, HMC and the mixture-MH (for suitable $\alpha\in[0,1]$) on a $10$-dimensional mixture of vMF distributions with $K=5$ vMF-mixture components on the unit sphere $\mathbb{S}^{9}$. We set $\kappa=100$ and draw $10^{6}$ samples.
@@
To gain some insight into the performance of the samplers on the multi-modal target, we first look at the marginal distributions (see Fig. 5). We estimate marginal histograms from the samples and compare them with the marginal distributions of the vMF mixture that we computed analytically. Both variants of geoSSS represent the underlying mixture significantly better compared to HMC and especially RWMH which seems to explore only a single component of the vMF mixture. The mixture-MH is almost on-par with our GeoSSS (shrink) variant. However, this represents the best result achieved by sweeping through $\alpha$ values [^1] For the $d=10,\kappa=100,K=5$ vMF mixture target, the best mixing probability $\alpha=0.2$ corresponds to the lowest KL divergence value (see Fig. 6B). For the systematic evaluation, we varied $\alpha\in[0,1]$ with a grid spacing of $0.1$. with $10^{6}$ MCMC steps for each case and therefore represents an additional computational burden of tuning this hyperparameter along with the step-size of RWMH for every target density.
Next we study quality measures to obtain more quantitative insights into the sampling performance. RWMH is the fastest to decorrelate, which is reflected by the rapid decay of the autocorrelation function (ACF) (see Fig. 6A). However, the reason for the fast decorrelation is that RWMH fails to explore the entire distribution, but is stuck in a single component. The geoSSS variants decorrelate much faster as compared to mixture-MH and HMC. Note that the slow, linear decay of the ACF for HMC results from the fact that it spends a large fraction of the simulation time in a single mode and only occasionally escapes to another mode (also HMC only visits three out of all five modes in the course of the simulation).
@@
where $\ell=(\ell_{1},\ldots,\ell_{K})$ with $\ell_{k}=1/K$ for $k\in\{1,\ldots,K\}$ is the ideal distribution and $q=(q_{1},\ldots,q_{K})\in[0,1]^{K}$. To estimate $q_{k}$, we assign each sample $x_{n}$ to the component under which $x_{n}$ is most likely, i.e.,
@@
Based on the allocations $k_{n}$, we can count how often each mode is visited by the Markov chain. A good MCMC sampler should produce $q_{k}\approx 1/K$ resulting in $\text{KL}(q\mid\ell)\approx 0$. The more the KL divergence differs from the minimum value of zero, the greater is the mismatch between the ideal and empirical frequency of mode visits. As seen from Fig. 6B, geoSSS (reject) achieves the lowest KL divergence whereas geoSSS (shrink) and mixture-MH perform comparably in achieving low KL divergence values. However, RWMH and HMC produce large KL divergences, indicating a complete failure to detect some of the components and, as a result, a poorer representation of the target distribution compared to geoSSS. Indeed, RWMH is stuck in a single mode in the course of the simulation. HMC fares better, but also misses two out of five modes.
@@
An efficient MCMC algorithm should explore the sphere rapidly by making large leaps from one sample to the next. The great circle distance between successive samples (see Fig. 6C) shows that the slice samplers and mixture-MH explore the sphere more efficiently than HMC and RWMH.
We further evaluate the performance of the samplers by fixing the components on the sphere and increasing the concentration parameter $\kappa$ from $50$ to $500$, therefore making the distributions “spikier”. We estimate the effective sample size (ESS) for each method by considering $10$ chains and $10^{6}$ MCMC steps from the first dimension, resulting in ESS values per method for each $\kappa$. Figure 7A shows the estimated relative [^2] Here and elsewhere the relative ESS refers to the estimated ESS divided by the total number of MCMC steps. ESS for each of the four samplers as a function of the concentration parameter $\kappa$. Overall, we observe that sampling the mixture model becomes more challenging as the components become more concentrated with increasing $\kappa$ for all methods. Notably, the geoSSS (reject) variant achieves the best relative ESS. We can again observe the trend that geoSSS (shrink) and mixture-MH (with the mixture hyperparameter fine-tuned for every $\kappa$ corresponding to the highest relative ESS) are comparable as well as that both outperform RWMH and HMC, particularly for lower $\kappa$ values. At large $\kappa$ values, all samplers perform similarly poorly and fail to produce a reliable approximation of the target.
@@
It is worth noting that although the rejection-based geoSSS yields the highest ESS, it is at the cost of generating a significantly larger number of rejected samples in comparison to the shrinkage-based geoSSS. This is demonstrated in Fig. 7B, where it is observed that the rejection-based geoSSS generates an increasingly larger number of rejected samples for higher values of $\kappa$. For instance, for $\kappa=50$, approximately 17 rejections happen per MCMC step, while for $\kappa=500$, around 60 rejections occur. On the other hand, the number of rejections for the shrinkage-based geoSSS increases at a much slower rate with $\kappa$. It can be seen here that for $\kappa=50$, we have approximately 4 rejections per MCMC step, and for $\kappa=500$, approximately just 6 rejections per MCMC step occur. The number of rejections directly determines the number of log probability evaluations and thereby the computational costs of the slice samplers. RWMH and HMC, on the other hand, have a fixed computational budget per MCMC step. In case of RWMH, one step requires a single log probability evaluation, whereas an HMC step involves one log probability evaluation as well as multiple gradient evaluations, one for each leapfrog step. In our settings, we use 10 leapfrog integration steps. In summary, RWMH has the lowest computational costs, but also the smallest ESS. The computational costs of shrinkage-based geoSSS is smaller than the costs of HMC, whilst it also performs better than HMC in terms of the other evaluation criteria.